Solve:


(π‘₯ βˆ’ 2𝑠𝑖𝑛𝑦 + 3)𝑑π‘₯ + (2π‘₯ βˆ’ 4𝑠𝑖𝑛𝑦 βˆ’ 3)π‘π‘œπ‘ π‘¦π‘‘π‘¦ = 0

1
Expert's answer
2022-02-16T17:09:57-0500

"\begin{gathered}
(x-2 \sin y+3) d x+(2 x-4 \sin y-3) \cos y d y=0 \\
u=\sin y \\
d u=\cos y d y \Rightarrow d y=d u / \cos y
\end{gathered}"

Substituting in differential equation.

Β "\begin{aligned}

&(x-2 u+3) d x+(2 x-4 u-3) \cos y \frac{d u}{\cos y}=0 \\

& \text { Integrating } \\

\Rightarrow & \int(x-2 u+3) d x+\int(2 x-4 u-3) d u=C \\

\Rightarrow & \frac{x^{2}}{2}-2u x+3 x+\frac{2 x^{2}}{2}-4u x-3 u=C \\

\Rightarrow & \frac{3 x^{2}-64 x-3 u+3 x=C}{2}\\\Rightarrow & \frac{3 x^{2}}{3}-6 x \sin y-3 \sin y+3 x=C

\end{aligned}"